A projectile is fired in such a way that its horizontal range is equal to its maximum height. What is the angle of projection?


1
Expert's answer
2021-11-07T19:28:37-0500
"R=3H""\frac{v^2\sin{2\theta}}{g}=2\frac{v^2\sin{\theta}\cos{\theta}}{g}=3
\frac{v^2\sin^2{\theta}}{2g}""\tan \theta=\frac{4}{3}""\theta=53\degree"